The electrical force on a 2-C charge is 20 N. What is the value of the electric field at the location of the charge?

a. 2 N/C
b.10 N/C
c.20 N/C
d. 40 N/C
e. 80 N/C
1
Expert's answer
2011-01-31T12:18:26-0500
The electric field can be written if such form:
E = F/q
Thus
E = 20/2 = 10 N/C
Answer: b.10 N/C